# MEDA: Dynamic KV Cache Allocation for Efficient Multimodal Long-Context Inference

### v1

@@ 1 Introduction
Long-context Multimodal Large Language Models (MLLMs) have achieved remarkable progress in processing multimodal long context involving long text-image and text-video inputs, as exemplified by LLaVA-NeXT Liu et al. (2024b), GPT-4V (Achiam et al., 2023) and long-video MLLMs Xue et al. (2024); Zhang et al. (2024b). These models are capable of handling complex multimodal patterns within their Key-Value (KV) caches, such as text accompanied by multiple interrelated images or lengthy video sequences, introducing intricate cross-modal interactions, as shown in Figure 1. However, despite these advancements, long-context MLLMs demand substantial resources as their multimodal KV caches grow with increasing input lengths in long-context settings, causing significant slowdown during inference.
@@
Conventional KV cache methods designed for text-only LLMs are difficult to be directly adopted to long-context multimodal inputs because they do not account for the complex cross-modal interactions present in long-context settings. Previous methods for KV cache compression in both text-only LLMs and MLLMs such as text-centric eviction-based methods (Zhang et al., 2023; Ren and Zhu, 2024; Li et al., 2024a), static progressive layer-wise reduction methods Zhang et al. (2024c); Yang et al. (2024), and multimodal pruning methods Wan et al. (2024b) have predominantly employed uniform cache size allocation across layers. However, these methods overlook the variations in attention density across different layers as illustrated in Figure 2. As a consequence, allocating a uniform KV cache size across different layers, without accounting for these variations, can not only lead to information loss in dense layers if less KV caches are allocated, resulting in reduced precision and suboptimal performance, but also cause significant inefficiency in sparse layers when more than enough KV caches are allocated.
@@
We evaluate MEDA across various recent MLLM backbones, including LLaVA-v1.5-13B Liu et al. (2023), LLaVA-NeXT-7B Liu et al. (2024b), and InternVL-v1.5-7B Chen et al. (2023) for multi-images tasks, as well as LLaVA-Video-7B/32B Zhang et al. (2024d), LongVA-7B Zhang et al. (2024b), and LongVILA-8B Xue et al. (2024) for long-video tasks. We also evaluate MEDA on diverse mutlimodal long-context datasets including MileBench (Song et al., 2024), Video-ChatGPT Maaz et al. (2023), DREAM-1K Wang et al. (2024a), and WorldQA Zhang et al. (2024e). Our results show that MEDA outperforms both state-of-the-art text-based and multimodal KV cache methods including H2O Zhang et al. (2024g), SnapKV Li et al. (2024a), PyramidKV Zhang et al. (2024c), LOOK-M Wan et al. (2024b), and is able to achieve up to 2.82 times faster inference speed and reduce KV cache memory footprint by up to 72%, while maintaining or improving performance on the target tasks.
Figure 2: Using the cross-modal attention entropy from Eq. 6, we analyze LLaVA-NeXT-7B Liu et al. (2024b) across different sub-tasks Song et al. (2024). We observe varying multimodal interaction patterns: early layers (e.g., Layer 1) exhibit dense attention weights with higher entropy, while deeper layers (e.g., Layer 24) exhibit sparse attention weights with lower entropy, given that they focus on key tokens (red columns), similar to the blue areas and text in Figure 1.
@@ 2 Related Work
Long-context MLLMs. Recent works have expanded MLLMs’ multimodal long-context capabilities through additional training. Liu et al. (2024a) leverage Blockwise RingAttention for scalable long-sequence training. LongVA Zhang et al. (2024b) first pre-trains LLMs on long-text sequences and then aligns Long LLMs using short vision data to generalize to multimodal long-text contexts. LongLLaVA Wang et al. (2024b) modifies the model architecture by integrating Mamba and Transformer blocks and employs a progressive training strategy using multiple images. Video-XL Shu et al. (2024) introduces visual context latent summarization to train models for handling even longer multimodal token sequences. In contrast, MEDA introduces a dynamic KV cache optimization algorithm, enhancing long-context multimodal inference without additional training and is compatible with these methods.
@@ 3.3 Multimodal KV Pair Selection and Merging
Multimodal KV Pair Selection. Text tokens often contain crucial semantic information in multimodal contexts Wan et al. (2024b). Therefore, during the prompt encoding stage, we prioritize the retention of text-based KV pairs and select conserved tokens based on the compression ratio and accumulated attention ranks. The tokens not selected are considered as less important tokens. Unlike previous eviction-based Zhang et al. (2024g) or static layer-wise reduction methods Zhang et al. (2024c); Yang et al. (2024) that discard these less important tokens, we employ KV cache merging techniques to preserve the integrity of the contextual information. The cumulative attention score $\mathbf{A}_{s}$ is computed as:
@@
Multimodal KV Pair Merging. To preserve the integrity of the contextual information, we employ a KV pair merging strategy that integrates the less important tokens into the conserved cache, rather than discarding them as in prior works Zhang et al. (2024c); Li et al. (2024a). Specifically, we perform a many-to-one nearest-neighbor matching Wan et al. (2024b) between the less important KV pairs ($\mathbf{K}_{less},\mathbf{V}_{less}$) and the conserved KV pairs ($\mathbf{K}_{c},\mathbf{V}_{c}$). We compute the similarity matrix $\mathbf{U}\in\mathbb{R}^{L_{\operatorname{les}}\times L_{c}}$ between the keys of less important and conserved tokens using cosine similarity:
@@ 4.2 Models
To evaluate MEDA’s robustness across different model architectures, we conduct experiments on both multimodal long-context and long-video models. For multimodal long-context models, we evaluate LLaVA-v1.5-13B Liu et al. (2023), LLaVA-NeXT-7B Liu et al. (2024b), and InternVL-v1.5-7B Chen et al. (2023). For long-video models, we evaluate LLaVA-Video-7B/32B Zhang et al. (2024d), LongVA-7B Zhang et al. (2024b), and LongVILA-8B Xue et al. (2024).
@@ 4.3 Baselines
To demonstrate the advantages of MEDA, we use the latest KV cache compression methods as baselines. H2O Zhang et al. (2024g) and SnapKV Li et al. (2024a) employ eviction-based strategies, while PyramidKV Zhang et al. (2024c) uses static layer-wise reduction. As these methods are text-centric, we adapt them for fair comparison in multimodal long-context settings. We also include LOOK-M Wan et al. (2024b), the state-of-the-art multimodal KV cache compression method.
@@ 8 Acknowledgment
This study is supported in part by NSF Award NeTS2312674.
@@ A.1.1 Details of MileBench
The MileBench dataset, introduced by (Song et al., 2024), is the first comprehensive benchmark specifically designed to evaluate the Multimodal Long-context capabilities of Multimodal Large Language Models (MLLMs). It contains 6,440 multimodal long-text samples sourced from 21 datasets, averaging 15.2 images and 422.3 words per sample. The benchmark is divided into two main components: Realistic Evaluation and Diagnostic Evaluation.

### v2

@@ 1 Introduction
Long-context Multimodal Large Language Models (MLLMs) have achieved remarkable progress in processing multimodal long context involving long text-image and text-video inputs, as exemplified by LLaVA-NeXT Liu et al. (2024c), GPT-4V (Achiam et al., 2023) and long-video MLLMs Xue et al. (2024); Zhang et al. (2024b). These models are capable of handling complex multimodal patterns within their Key-Value (KV) caches, such as text accompanied by multiple interrelated images or lengthy video sequences, introducing intricate cross-modal interactions, as shown in Figure 1. However, despite these advancements, long-context MLLMs demand substantial resources as their multimodal KV caches grow with increasing input lengths in long-context settings, causing significant slowdown during inference.
@@
Conventional KV cache methods designed for text-only LLMs are difficult to be directly adopted to long-context multimodal inputs because they do not account for the complex cross-modal interactions present in long-context settings. Previous methods for KV cache compression in both text-only LLMs and MLLMs such as text-centric eviction-based methods (Zhang et al., 2023; Ren and Zhu, 2024; Li et al., 2024a), static progressive layer-wise reduction methods Zhang et al. (2024c); Yang et al. (2024), and multimodal pruning methods Wan et al. (2024c) have predominantly employed uniform cache size allocation across layers. However, these methods overlook the variations in attention density across different layers as illustrated in Figure 2. As a consequence, allocating a uniform KV cache size across different layers, without accounting for these variations, can not only lead to information loss in dense layers if less KV caches are allocated, resulting in reduced precision and suboptimal performance, but also cause significant inefficiency in sparse layers when more than enough KV caches are allocated.
@@
We evaluate MEDA across various recent MLLM backbones, including LLaVA-v1.5-13B Liu et al. (2023), LLaVA-NeXT-7B Liu et al. (2024c), and InternVL-v1.5-7B Chen et al. (2023) for multi-images tasks, as well as LLaVA-Video-7B/32B Zhang et al. (2024d), LongVA-7B Zhang et al. (2024b), and LongVILA-8B Xue et al. (2024) for long-video tasks. We also evaluate MEDA on diverse mutlimodal long-context datasets including MileBench (Song et al., 2024), Video-ChatGPT Maaz et al. (2023), DREAM-1K Wang et al. (2024a), and WorldQA Zhang et al. (2024e). Our results show that MEDA outperforms both state-of-the-art text-based and multimodal KV cache methods including H2O Zhang et al. (2024g), SnapKV Li et al. (2024a), PyramidKV Zhang et al. (2024c), LOOK-M Wan et al. (2024c), and is able to achieve up to 2.82 times faster inference speed and reduce KV cache memory footprint by up to 72%, while maintaining or improving performance on the target tasks.
Figure 2: Using the cross-modal attention entropy from Eq. 6, we analyze LLaVA-NeXT-7B Liu et al. (2024c) across different sub-tasks Song et al. (2024). We observe varying multimodal interaction patterns: early layers (e.g., Layer 1) exhibit dense attention weights with higher entropy, while deeper layers (e.g., Layer 24) exhibit sparse attention weights with lower entropy, given that they focus on key tokens (red columns), similar to the blue areas and text in Figure 1.
@@ 2 Related Work
Long-context MLLMs. Recent works have expanded MLLMs’ multimodal long-context capabilities through additional training. Liu et al. (2024b) leverage Blockwise RingAttention for scalable long-sequence training. LongVA Zhang et al. (2024b) first pre-trains LLMs on long-text sequences and then aligns Long LLMs using short vision data to generalize to multimodal long-text contexts. LongLLaVA Wang et al. (2024b) modifies the model architecture by integrating Mamba and Transformer blocks and employs a progressive training strategy using multiple images. Video-XL Shu et al. (2024) introduces visual context latent summarization to train models for handling even longer multimodal token sequences. In contrast, MEDA introduces a dynamic KV cache optimization algorithm, enhancing long-context multimodal inference without additional training and is compatible with these methods.
@@ 3.3 Multimodal KV Pair Selection and Merging
Multimodal KV Pair Selection. Text tokens often contain crucial semantic information in multimodal contexts Wan et al. (2024c). Therefore, during the prompt encoding stage, we prioritize the retention of text-based KV pairs and select conserved tokens based on the compression ratio and accumulated attention ranks. The tokens not selected are considered as less important tokens. Unlike previous eviction-based Zhang et al. (2024g) or static layer-wise reduction methods Zhang et al. (2024c); Yang et al. (2024) that discard these less important tokens, we employ KV cache merging techniques to preserve the integrity of the contextual information. The cumulative attention score $\mathbf{A}_{s}$ is computed as:
@@
Multimodal KV Pair Merging. To preserve the integrity of the contextual information, we employ a KV pair merging strategy that integrates the less important tokens into the conserved cache, rather than discarding them as in prior works Zhang et al. (2024c); Li et al. (2024a). Specifically, we perform a many-to-one nearest-neighbor matching Wan et al. (2024c) between the less important KV pairs ($\mathbf{K}_{less},\mathbf{V}_{less}$) and the conserved KV pairs ($\mathbf{K}_{c},\mathbf{V}_{c}$). We compute the similarity matrix $\mathbf{U}\in\mathbb{R}^{L_{\operatorname{les}}\times L_{c}}$ between the keys of less important and conserved tokens using cosine similarity:
@@ 4.2 Models
To evaluate MEDA’s robustness across different model architectures, we conduct experiments on both multimodal long-context and long-video models. For multimodal long-context models, we evaluate LLaVA-v1.5-13B Liu et al. (2023), LLaVA-NeXT-7B Liu et al. (2024c), and InternVL-v1.5-7B Chen et al. (2023). For long-video models, we evaluate LLaVA-Video-7B/32B Zhang et al. (2024d), LongVA-7B Zhang et al. (2024b), and LongVILA-8B Xue et al. (2024).
@@ 4.3 Baselines
To demonstrate the advantages of MEDA, we use the latest KV cache compression methods as baselines. H2O Zhang et al. (2024g) and SnapKV Li et al. (2024a) employ eviction-based strategies, while PyramidKV Zhang et al. (2024c) uses static layer-wise reduction. As these methods are text-centric, we adapt them for fair comparison in multimodal long-context settings. We also include LOOK-M Wan et al. (2024c), the state-of-the-art multimodal KV cache compression method.
@@ 8 Acknowledgment
This study is supported in part by NSF Award NeTS-2312675.
@@ A.1.1 Details of MileBench
The MileBench dataset, introduced by (Song et al., 2024), is the first comprehensive benchmark specifically designed to evaluate the Multimodal Long-context capabilities of Multimodal Large Language Models (MLLMs) Huang et al. (2024); Zhu et al. (2024); Gong et al. (2024); Liu et al. (2024a); Wan et al. (2024a), which is different from static classification tasks Wan (2023); Chen et al. (2025). It contains 6,440 multimodal long-text samples sourced from 21 datasets, averaging 15.2 images and 422.3 words per sample. The benchmark is divided into two main components: Realistic Evaluation and Diagnostic Evaluation.